# LanPaint: Training-Free Diffusion Inpainting with Asymptotically Exact and Fast Conditional Sampling

### v2

@@ Abstract
Diffusion models excel at joint pixel sampling for image generation but lack efficient training-free methods for partial conditional sampling (e.g., inpainting with known pixels). Prior work typically formulates this as an intractable inverse problem, relying on coarse variational approximations, heuristic losses requiring expensive backpropagation, or slow stochastic sampling. These limitations preclude: (1) accurate distributional matching in inpainting results, (2) efficient inference modes without gradient, (3) compatibility with fast ODE-based samplers. To address these limitations, we propose LanPaint: a training-free, asymptotically exact partial conditional sampling methods for ODE-based and rectified flow diffusion models. By leveraging carefully designed Langevin dynamics, LanPaint enables fast, backpropagation-free Monte Carlo sampling. Experiments demonstrate that our approach achieves superior performance with precise partial conditioning and visually coherent inpainting across diverse tasks.
@@
Figure 1: Demonstration of LanPaint-5 (5 inner iterations) inpainting results on HiDream-L1 (HiDream.ai, 2025), Flux.1 dev (Labs, 2024), SD 3.5 (Esser et al., 2024) and XL (Podell et al., 2023). Images are generated through ComfyUI (Comfy Org, 2025) with Euler sampler Karras et al. (2022) (30 steps). All samples generated from a fixed seed (seed=0) producing a batch of 4 distinct random latents to avoid cherry-picking. These results demonstrate LanPaint’s practical effectiveness across modern diffusion architectures, including both rectified flow (HiDream, Flux, SD 3.5) and denoising (SD XL) models.
@@ 1 Introduction
Denoising Diffusion Probabilistic Models (DDPM) Sohl-Dickstein et al. (2015); Song & Ermon (2019a); Song et al. (2020b); Ho et al. (2020); Rombach et al. (2021); Betker et al. (2023) have emerged as powerful generative frameworks that produce high-quality outputs through iterative denoising. Subsequent advances in ODE-based deterministic samplers Karras et al. (2022); Lu et al. (2022); Zhao et al. (2023), as well as equivalent rectified flow models (Lipman et al., 2022; Liu et al., 2022b; Gao et al., 2025) have dramatically improved DDPM’s efficiency, reducing sampling steps from hundreds to dozens. These innovations, combined with community-trained model variants, have broadened the scope and quality of generative visual arts.
@@
While diffusion models excel at whole-image sampling, their global denoising mechanism inherently limits partial conditional sampling given partially known pixels. Mathematically, for a pretrained model $p(\mathbf{z})$ with arbitrary splitting $\mathbf{z}=(\mathbf{x},\mathbf{y})$, sampling $\mathbf{x}\sim p(\mathbf{x}|\mathbf{y})$ in a training-free way remains a fundamental challenge for diffusion models. Current approaches fall into two categories: (1) Sequential Monte Carlo (SMC) methods Trippe et al. (2022); Wu et al. (2024). They depend on stochastic DDPM sampling, making them incompatible with deterministic ODE samplers and computationally expensive; and (2) Langevin Dynamics Monte Carlo (LMC) methods Lugmayr et al. (2022); Cornwall et al. (2024). They can be treated as iteratively denoise and renoise hence are compatible with ODE samplers. But they suffer from convergence issues Cornwall et al. (2024) and local maxima trapping - a key limitation we analyze in this work.
@@
Another line of work formulates inpainting as a linear inverse problem, where observed pixels $\mathbf{y}=H\mathbf{z}+\epsilon$ arise from a known degenerate operator $H$ and Gaussian noise $\epsilon$. These methods approximate the intractable posterior $q(\mathbf{z}|\mathbf{y})$ using the diffusion prior $p(\mathbf{z})$ through either heuristic losses ($\|\mathbf{y}-H\mathbf{z}\|_{2}^{2}$) or DDIM-based variational inference (Chung et al., 2022a; b; Grechka et al., 2024; Kawar et al., 2022; Zhang et al., 2023a; Janati et al., 2024; Ben-Hamu et al., 2024). While linear inverse problem is applicable to other generative models (e.g., GANs (Goodfellow et al., 2014)) and tasks like deblurring, they fundamentally address a different problem: The marginal $q(\mathbf{x}|\mathbf{y})$ of posterior $q(\mathbf{z}|\mathbf{y})$ is a heuristic approximation due to its intractable nature, rather than being constrained to match the joint distribution $p(\mathbf{z})=p(\mathbf{x},\mathbf{y})$.
@@
Training-based approaches (Zhang et al., 2023b; Mayet et al., 2024; Zhuang et al., 2024) also address conditional sampling and achieve good performance. However, these methods require training specialized networks or modules for each model architecture, making them impractical for adoption across business and community models of diverse architecture, hindering their ecosystem development.
@@
In this work, we propose LanPaint, a training-free and efficient partial conditional sampling method based on Langevin Dynamics Monte Carlo, tailored for ODE-based diffusion samplers and rectified flow models. LanPaint achieves asymptotically exact partial conditional sampling without heuristics. It introduces two core innovations: (1) Bidirectional Guided (BiG) Score, which enables mutual adaptation between inpainted and observed regions, avoids local maxima traps caused by ODE-samplers with large diffusion step sizes. This significantly improves in-painting quality; and (2) Fast Langevin Dynamics (FLD), an accelerated Langevin sampling scheme that yields high-fidelity results in just 5 inner iterations per step, drastically reducing computational costs compared to prior Langevin methods. Experiments confirm that LanPaint outperforms existing training-free approaches, delivering high-quality inpainting and outpainting results for both pixel-space and latent-space models.
@@ 2.1 Langevin Dynamics
Langevin Dynamics is a Monte Carlo sampling techniques. For a target distribution $p(\mathbf{z})$, the dynamics is governed by the stochastic differential equation (SDE):
@@
where $\mathbf{s}(\mathbf{z})=\nabla_{\mathbf{z}}\log p(\mathbf{z})$. It asymptotically converges to the stationary distribution $\mathbf{z}_{\tau}\sim p(\mathbf{z})$ as $\tau\to\infty$ (Appendix A). However, this method risks trapping samples at local likelihood maximas of $p(\mathbf{z})$.
@@ 2.2 DDPM and ODE Based Sampling
where $\bar{\alpha}_{t}=e^{-t}$ and $d\mathbf{W}$ is a Brownian motion increment (Appendix B). The OU process ensures $\mathbf{z}_{t}$ transitions smoothly from the data distribution $p(\mathbf{z})$ to pure noise $\mathcal{N}(\mathbf{0},\mathbf{I})$.
@@
[…]
where $t^{\prime}\in[0,\,T]$ is the backward time, $t=T-t^{\prime}$, and $\mathbf{s}(\mathbf{z},t)=\nabla_{\mathbf{z}}\log p_{t}(\mathbf{z})$ is the score function of the random variable $\mathbf{z}_{t}$, which guides noise removal. The score function is usually learnt as a denoising neural network (Appendix B, C). The recent popular flow matching models, though commonly thought as a different architecture, also fails into this category (Appendix D).
@@ 2.3 InPainting as Partial Conditional Sampling
Unlike many works treating inpainting as solving an ill-posed inverse problem, we treat inpainting as a partial conditional sampling problem for diffusion models: given a joint distribution of images (DDPM), how to sample one part given the other part of the image. Partial conditional sampling in DDPMs poses a significant challenge due to the inaccessibility of the conditional score function. A DDPM is trained to model a joint distribution $p(\mathbf{z})=p(\mathbf{x},\mathbf{y})$. But partial conditional sampling aims to generate $\mathbf{x}\sim p(\mathbf{x}\mid\mathbf{y}=\mathbf{y}_{o})$. During sampling process, DDPM’s denoising network is trained to provide these joint scores:
@@
$\mathbf{s}_{\mathbf{x}|\mathbf{y}_{0}}=\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x}\mid\mathbf{y}_{0})=\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x},\mathbf{y}\mid\mathbf{y}_{0}),$ (6)
which is required for direct sampling from $p(\mathbf{x}\mid\mathbf{y}_{0})$, remains inaccessible. (The second equality holds because $\mathbf{x}$ and $\mathbf{y}$ are conditionally independent given $\mathbf{y}_{0}$.)
Decoupling Approximation Rather than tracking the unknown distribution $p_{t}(\mathbf{x},\mathbf{y}\mid\mathbf{y}_{0})$, we can approximate it as an alternative distribution $q_{t}(\mathbf{x},\mathbf{y}\mid\mathbf{y}_{0})$:
$p_{t}(\mathbf{x},\mathbf{y}\mid\mathbf{y}_{0})\approx q_{t}(\mathbf{x},\mathbf{y}\mid\mathbf{y}_{0})=p_{t}(\mathbf{x}\mid\mathbf{y})\cdot p_{t}(\mathbf{y}\mid\mathbf{y}_{0}).$ (7)
This decoupling introduces dependencies between $\mathbf{x}_{t}$ and $\mathbf{y}_{t}$—unlike the original DDPM framework, where $\mathbf{x}_{t}$ and $\mathbf{y}_{t}$ remain independent given $\mathbf{y}_{o}$. However, the approximation ($\approx$) becomes exact ($=$) at $t=0$ because $p_{t=0}(\mathbf{y}\mid\mathbf{y}_{0})=\delta(\mathbf{y}-\mathbf{y}_{0})$, ensuring the final output precisely follows $p(\mathbf{x}\mid\mathbf{y}=\mathbf{y}_{0})$.
Here, $p_{t}(\mathbf{y}\mid\mathbf{y}_{o})$ is analytically known from the forward process $p_{t}(\mathbf{y}\mid\mathbf{y}_{o})=\mathcal{N}\!\bigl{(}\mathbf{y}\mid\sqrt{\bar{\alpha}_{t}}\mathbf{y}_{o},(1-\bar{\alpha}_{t})\mathbf{I}\bigr{)},$ while $p_{t}(\mathbf{x}\mid\mathbf{y})$ shares the same score $\mathbf{s}_{\mathbf{x}}$ as the joint distribution in equation 5. This makes the approximation tractable in practice. The only problem remains is how to let the DDPM generate samples from $q_{t}(\mathbf{x},\mathbf{y}\mid\mathbf{y}_{0})$ instead of $p_{t}(\mathbf{x},\mathbf{y})$ during the sampling process.
The Replace Method During backward sampling with $t^{\prime}=T-t$, Song & Ermon (2019b) approximately samples $\mathbf{x}_{t^{\prime}},\mathbf{y}_{t^{\prime}}\sim q_{T-t^{\prime}}(\mathbf{x},\mathbf{y}\mid\mathbf{y}_{0})$ by replacing unconditionally sampled $\mathbf{y}_{t^{\prime}}$ with
@@
for each time step of a sampling process. It correctly samples $\mathbf{y}_{t^{\prime}}$, but fail to ensure $\mathbf{x}_{t^{\prime}}\sim p_{T-t^{\prime}}(\mathbf{x}|\mathbf{y})$. This yields a mix between unconditional and partial conditional sampling
$(\mathbf{x}_{t^{\prime}},\mathbf{y}_{t^{\prime}})\sim\text{Between}\,[\,p_{T-t^{\prime}}(\mathbf{x},\mathbf{y})\,]\;\text{and}\;[\,q_{T-t^{\prime}}(\mathbf{x},\mathbf{y}\mid\mathbf{y}_{0})\,].$ (9)
@@
RePaint (Lugmayr et al., 2022) improves over the Replace Method by introducing a “time traveling” step that refines $\mathbf{x}_{t^{\prime}}$. Between two backward diffusion time $t^{\prime}_{i}=T-t_{i}$ and $t^{\prime}_{i+1}=T-t_{i+1}$, it performs multiple inner iterations, alternating between forward and backward diffusion steps:
@@
Meanwhile, $\mathbf{y}_{t^{\prime}}$ is still replaced at each step by equation 8. As this work focusing on ODE sampling, we replace RePaint’s SDE backward steps with an ODE backward (Euler ODE sampler Karras et al. (2022)), and name this adaptation Repaint-Euler.
By adding forward and backward together, Repaint effectively simulates a Langevin dynamics equation 1 with $d\tau=t^{\prime}_{i+1}-t^{\prime}_{i}$, whose stationary distribution is $p_{t_{i}}(\mathbf{x}\mid\mathbf{y})$. Therefore after sufficient iterations, RePaint asymptotically produce samples from $q_{t}(\mathbf{x},\mathbf{y}\mid\mathbf{y}_{0})$.
RePaint’s key limitation is that its step size $d\tau$ is fixed by the backward sampling schedule. A small number of sampling steps leads to overly large steps, causing divergence, while too many steps result in excessively small steps, preventing convergence to a stationary state.

### v3

@@ Abstract
Diffusion models excel at joint pixel sampling for image generation but lack efficient training-free methods for partial conditional sampling (e.g., inpainting with known pixels). Prior works typically formulate this as an intractable inverse problem, relying on coarse variational approximations, heuristic losses requiring expensive backpropagation, or slow stochastic sampling. These limitations preclude (1) accurate distributional matching in inpainting results, (2) efficient inference modes without gradients, and (3) compatibility with fast ODE-based samplers. To address these limitations, we propose LanPaint: a training-free, asymptotically exact partial conditional sampling method for ODE-based and rectified-flow diffusion models. By leveraging carefully designed Langevin dynamics, LanPaint enables fast, backpropagation-free Monte Carlo sampling. Experiments demonstrate that our approach achieves superior performance with precise partial conditioning and visually coherent inpainting across diverse tasks. Code is available on https://github.com/scraed/LanPaint.
@@
Figure 1: Demonstration of LanPaint-5 (5 inner iterations) inpainting results on HiDream-L1 (HiDream.ai, 2025), Flux.1 dev (Labs, 2024), SD 3.5 (Esser et al., 2024) and XL (Podell et al., 2023). Images are generated through ComfyUI (Comfy Org, 2025) with Euler sampler (Karras et al., 2022) (30 steps). All samples generated from a fixed seed (seed=0) producing a batch of 4 distinct random latents to avoid cherry-picking. These results demonstrate LanPaint’s practical effectiveness across modern diffusion architectures, including both rectified flow (HiDream, Flux, SD 3.5) and denoising (SD XL) models.
@@ 1 Introduction
Denoising Diffusion Probabilistic Models (DDPMs) (Sohl-Dickstein et al., 2015; Song & Ermon, 2019a; Song et al., 2020c; Ho et al., 2020; Rombach et al., 2021; Betker et al., 2023) have emerged as powerful generative frameworks that produce high-quality outputs through iterative denoising. Subsequent advances in ODE-based deterministic samplers (Karras et al., 2022; Lu et al., 2022; Zhao et al., 2023), as well as equivalent rectified flow models (Lipman et al., 2022; Liu et al., 2022b; Gao et al., 2025) have dramatically improved the’ efficiency of DDPMs, reducing the sampling steps from hundreds to dozens. These innovations, combined with model variants trained in the community, have broadened the scope and quality of the generative visual arts.
While diffusion models excel at whole-image sampling, their global denoising mechanism inherently limits partial conditional sampling given partially known pixels. The key question is
Given $p(\mathbf{x},\mathbf{y})$, how to sample from $p(\mathbf{x}|\mathbf{y})$?
@@
More rigorously, for a pretrained model $p(\mathbf{z})$ with arbitrary splitting $\mathbf{z}=(\mathbf{x},\mathbf{y})$, sampling $\mathbf{x}\sim p(\mathbf{x}|\mathbf{y})$ in a training-free way remains a fundamental challenge for diffusion models. Current approaches fall into two categories: (1) Sequential Monte Carlo (SMC) methods (Trippe et al., 2022; Wu et al., 2024). They depend on stochastic DDPM sampling, making them incompatible with deterministic ODE samplers and computationally expensive; and (2) Langevin Dynamics Monte Carlo (LMC) methods (Lugmayr et al., 2022; Cornwall et al., 2024). They can be treated as iterative denoising and renoising, making them compatible with ODE samplers. But they suffer from convergence issues (Cornwall et al., 2024) and local maxima trapping - a key limitation we analyze in this work.
@@
Another line of work formulates inpainting as a linear inverse problem, where observed pixels $\mathbf{y}=H\mathbf{z}+\epsilon$ arise from a known degenerate operator $H$ and Gaussian noise $\epsilon$. These methods approximate the intractable posterior $q(\mathbf{z}|\mathbf{y})$ using the diffusion prior $p(\mathbf{z})$ through either heuristic losses ($\|\mathbf{y}-H\mathbf{z}\|_{2}^{2}$) or DDIM-based variational inference (Chung et al., 2022a; b; Grechka et al., 2024; Kawar et al., 2022; Zhang et al., 2023a; Janati et al., 2024; Ben-Hamu et al., 2024). While linear inverse problem is applicable to other generative models (e.g., GANs (Goodfellow et al., 2014)) and tasks such as deblurring, they fundamentally address a different problem: The posterior $q(\mathbf{z}|\mathbf{y})$ is a heuristic approximation that aims to construct a visually plausible $\mathbf{z}=(\mathbf{x},\mathbf{y})$ without requiring $\mathbf{z}$ to follow exactly the joint distribution $p(\mathbf{z})$ modeled by the pretrained diffusion model.
@@
Training-based approaches (Zhang et al., 2023b; Mayet et al., 2024; Zhuang et al., 2024) also address conditional sampling and achieve good performance. However, these methods require training of specialized networks or modules for each model architecture, making them impractical for adoption across business and community models of diverse architectures, hindering their ecosystem development.
@@
In this work, we propose LanPaint, a training-free and efficient partial conditional sampling method based on Langevin Dynamics Monte Carlo, tailored for ODE-based diffusion samplers and rectified flow models. LanPaint achieves asymptotically exact partial conditional sampling without heuristics. It introduces two core innovations: (1) Bidirectional Guided (BiG) Score, which enables mutual adaptation between inpainted and observed regions, and avoids local maxima traps caused by ODE-samplers with large diffusion step sizes. This significantly improves inpainting quality; and (2) Fast Langevin Dynamics (FLD), an accelerated Langevin sampling scheme that yields high-fidelity results in just 5 inner iterations per step, drastically reducing computational costs compared to prior Langevin methods. Experiments confirm that LanPaint outperforms existing training-free approaches, delivering high-quality inpainting and outpainting results for both pixel-space and latent-space models.
2 Related Works
2.1 ODE-based Sampling Methods and Rectified Flow
Vanilla DDPMs are slow, requiring numerous denoising steps. Acceleration strategies like approximate diffusion processes (Song et al., 2020c; Liu et al., 2022a; Song et al., 2020a; Zhao et al., 2023) and advanced ODE solvers (Karras et al., 2022; Lu et al., 2022; Zhao et al., 2023) convert stochastic DDPM sampling into deterministic ODE flows, enabling larger time steps and faster generation, dominating current diffusion model sampling. Rectified flow (Lipman et al., 2022; Liu et al., 2022b), a recent alternative, is a reparameterization of ODE-based diffusion models with improved numerical properties. As shown by (Gao et al., 2025), it also belongs to the ODE sampling family.
LMC
One family of works tackling this problem is Langevin dynamics Monte Carlo (LMC). This approach was pioneered by RePaint (Lugmayr et al., 2022), which employs a "time travel" mechanism of iterative denoising and renoising steps. This mechanism was later shown by (Cornwall et al., 2024) to be equivalent to LMC. A crucial advantage of this formulation is that it enables easy switching between stochastic and ODE sampling, by simply switching the denoising step from SDE to ODE.
In this work, we first adapt both RePaint and the Langevin dynamics approach from (Cornwall et al., 2024) to an ODE sampler as our baseline. Through this implementation, we identify their common limitation—susceptibility to local maxima trapping—and subsequently address it with our proposed bidirectional guidance.
SMC
Linear Inverse Problems
One approach minimizes heuristic losses $\|\mathbf{y}-H\mathbf{z}\|_{2}^{2}$, as in MCG (Chung et al., 2022b), DPS (Chung et al., 2022a), GradPaint (Grechka et al., 2024), DCPS (Janati et al., 2024), and D-Flow (Ben-Hamu et al., 2024). However, most of these methods are tailored for stochastic DDPM samplers without easy migration to ODE samplers (e.g. DCPS), requiring costly full-model differentiation or expensive optimization (e.g., line search in D-Flow). For our baselines, we select only those compatible with ODE samplers and free from line search.
Another approach uses variational inference in the DDIM framework, such as DDRM (Kawar et al., 2022). CoPaint (Zhang et al., 2023a) and MMPS (Rozet et al., 2024) also adopt variational inference and expectation-maximization to improve the optimization of heuristic losses and enhance stability. While DDIM enables fast deterministic sampling, its variational approximations introduce limitations.
While we include these inverse problem baselines for comparison, they differ fundamentally from partial conditional sampling. Linear inverse problems (effective for inpainting) do not enforce matching the joint distribution between inpainted and known regions—a core requirement of our approach. Moreover, minimizing the heuristic loss typically requires 2–4 times more GPU memory than standard inference, making it prohibitive for production-level models on consumer GPUs. Accordingly, these baselines are included as supplementary reference points rather than essential benchmarks.
2.3 Trained Partial Conditional Sampling with Diffusion Models
@@ 3.1 Langevin Dynamics
Langevin Dynamics is a Monte Carlo sampling technique. For a target distribution $p(\mathbf{z})$, the dynamics is governed by the stochastic differential equation (SDE):
@@
where $\mathbf{s}(\mathbf{z})=\nabla_{\mathbf{z}}\log p(\mathbf{z})$. It asymptotically converges to the stationary distribution $\mathbf{z}_{\tau}\sim p(\mathbf{z})$ as $\tau\to\infty$ (Appendix B). However, this method risks trapping samples at local likelihood maxima of $p(\mathbf{z})$.
@@ 3.2 DDPM and ODE Based Sampling
where $\bar{\alpha}_{t}=e^{-t}$ and $d\mathbf{W}$ is a Brownian motion increment (Appendix C). The OU process ensures $\mathbf{z}_{t}$ transitions smoothly from the data distribution $p(\mathbf{z})$ to pure noise $\mathcal{N}(\mathbf{0},\mathbf{I})$.
@@
$\text{SDE: }d\mathbf{z}_{t^{\prime}}=\left(\tfrac{1}{2}\mathbf{z}_{t^{\prime}}+\mathbf{s}(\mathbf{z}_{t^{\prime}},T-t^{\prime})\right)dt^{\prime}+d\mathbf{W}_{t^{\prime}};\quad\text{ODE: }d\mathbf{z}_{t^{\prime}}=\tfrac{1}{2}\left(\mathbf{z}_{t^{\prime}}+\mathbf{s}(\mathbf{z},\,T-t^{\prime})\right)dt^{\prime}.$ (4)
where $t^{\prime}\in[0,\,T]$ is the backward time, $t=T-t^{\prime}$, and $\mathbf{s}(\mathbf{z},t)=\nabla_{\mathbf{z}}\log p_{t}(\mathbf{z})$ is the score function of the random variable $\mathbf{z}_{t}$, which guides noise removal. The score function is usually learnt as a denoising neural network (Appendix C and D). The recent popular flow matching model, though commonly thought as a different architecture, also falls into this category (Appendix E).
@@ 3.3 Inpainting as Partial Conditional Sampling
Unlike many works treating inpainting as solving an ill-posed inverse problem, we treat inpainting as a partial conditional sampling problem for diffusion models: given a joint distribution of images (DDPM), how to sample one part given the other part of the image. Partial conditional sampling in DDPMs poses a significant challenge due to the inaccessibility of the conditional score function. A DDPM is trained to model a joint distribution $p(\mathbf{z})=p(\mathbf{x},\mathbf{y})$, where $\mathbf{z}$ is the whole image, $\mathbf{x}$ denotes the region to be inpainted, and $\mathbf{y}$ denotes the region to be observed. But partial conditional sampling aims to generate $\mathbf{x}\sim p(\mathbf{x}\mid\mathbf{y}=\mathbf{y}_{o})$, where $\mathbf{y}_{o}$ is the observed region of a given image. During the sampling process, DDPM’s denoising network is trained to provide these joint scores:
@@
$\mathbf{s}_{\mathbf{x}|\mathbf{y}_{o}}=\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x}\mid\mathbf{y}_{o})=\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x},\mathbf{y}\mid\mathbf{y}_{o}),$ (6)
which is required for direct sampling from $p(\mathbf{x}\mid\mathbf{y}_{o})$, remains inaccessible. (The second equality holds because $\mathbf{x}$ and $\mathbf{y}$ are conditionally independent given $\mathbf{y}_{o}$.)
Decoupling Approximation Rather than tracking the unknown distribution $p_{t}(\mathbf{x},\mathbf{y}\mid\mathbf{y}_{o})$, we can approximate it as an alternative distribution $q_{t}(\mathbf{x},\mathbf{y}\mid\mathbf{y}_{o})$:
$p_{t}(\mathbf{x},\mathbf{y}\mid\mathbf{y}_{o})\approx q_{t}(\mathbf{x},\mathbf{y}\mid\mathbf{y}_{o})=p_{t}(\mathbf{x}\mid\mathbf{y})\cdot p_{t}(\mathbf{y}\mid\mathbf{y}_{o}).$ (7)
This decoupling introduces dependencies between $\mathbf{x}_{t}$ and $\mathbf{y}_{t}$ - unlike the original DDPM framework, where $\mathbf{x}_{t}$ and $\mathbf{y}_{t}$ are independent given $\mathbf{y}_{o}$. In particular, the approximation ($\approx$) becomes exact (==) at $t=0$ because $p_{t=0}(\mathbf{y}\mid\mathbf{y}_{o})=\delta(\mathbf{y}-\mathbf{y}_{o})$, ensuring that the final output still follows precisely $p(\mathbf{x}\mid\mathbf{y}=\mathbf{y}_{o})$.
Here, $p_{t}(\mathbf{y}\mid\mathbf{y}_{o})$ is analytically known from the forward process $p_{t}(\mathbf{y}\mid\mathbf{y}_{o})=\mathcal{N}\!\bigl(\mathbf{y}\mid\sqrt{\bar{\alpha}_{t}}\mathbf{y}_{o},(1-\bar{\alpha}_{t})\mathbf{I}\bigr),$ while $p_{t}(\mathbf{x}\mid\mathbf{y})$ shares the same score $\mathbf{s}_{\mathbf{x}}$ as the joint distribution in Eq.5. This makes the approximation tractable in practice. The only problem that remains is how to let the DDPM generate samples from $q_{t}(\mathbf{x},\mathbf{y}\mid\mathbf{y}_{o})$ instead of $p_{t}(\mathbf{x},\mathbf{y})$ during the sampling process.
The Replace Method During backward diffusion sampling with $t^{\prime}=T-t$, (Song & Ermon, 2019b) approximately samples $\mathbf{x}_{t^{\prime}},\mathbf{y}_{t^{\prime}}\sim q_{T-t^{\prime}}(\mathbf{x},\mathbf{y}\mid\mathbf{y}_{o})$ by replacing unconditionally sampled $\mathbf{y}_{t^{\prime}}$ with
@@
for each time step of a sampling process. It correctly samples $\mathbf{y}_{t^{\prime}}$, but fails to ensure $\mathbf{x}_{t^{\prime}}\sim p_{T-t^{\prime}}(\mathbf{x}|\mathbf{y})$. This yields a mix between unconditional and partial conditional sampling
$(\mathbf{x}_{t^{\prime}},\mathbf{y}_{t^{\prime}})\sim\text{Between}\,[\,p_{T-t^{\prime}}(\mathbf{x},\mathbf{y})\,]\;\text{and}\;[\,q_{T-t^{\prime}}(\mathbf{x},\mathbf{y}\mid\mathbf{y}_{o})\,].$ (9)
@@
RePaint (Lugmayr et al., 2022) improves over the Replace Method by introducing a “time traveling” step that refines $\mathbf{x}_{t^{\prime}}$. Between two backward diffusion times $t^{\prime}_{i}=T-t_{i}$ and $t^{\prime}_{i+1}=T-t_{i+1}$, it performs multiple inner iterations, alternating between forward and backward diffusion steps:
@@
Meanwhile, $\mathbf{y}_{t^{\prime}}$ is still replaced at each step by Eq.8. As this work focuses on ODE sampling, we replace RePaint’s SDE backward steps with an ODE backward (Euler ODE sampler (Karras et al., 2022)), and name this adaptation Repaint-Euler.
By adding forward and backward together, Repaint effectively simulates a Langevin dynamics Eq.1 with $d\tau=t^{\prime}_{i+1}-t^{\prime}_{i}$, whose stationary distribution is $p_{t_{i}}(\mathbf{x}\mid\mathbf{y})$. Therefore after sufficient iterations, RePaint asymptotically produces samples from $q_{t}(\mathbf{x},\mathbf{y}\mid\mathbf{y}_{o})$.
RePaint’s key limitation is that its step size $d\tau$ is fixed by the backward sampling schedule. A small number of sampling steps leads to overly large step sizes, causing divergence, while too many steps result in excessively small step sizes, preventing convergence to a stationary state.